a car travelling at 60kph left town at 9am. another car travelling at 65kph left the same place half an hour later to overtake the first. at what time did the second car overtake the first?
1
Expert's answer
2018-11-26T16:12:13-0500

Answer on Question #83265

a car travelling at 60kph left town at 9am. another car travelling at 65kph left the same place half an hour later to overtake the first. at what time did the second car overtake the first?

Solution

The first car will travel distance S1= t1*60

The second car will travel the same distance S1= t2*65, where t2= t1 -0.5

S1 =S1

t1*60 = (t1- 0.5)*65

t1= 6.5

As the first car started at 9 am, then the time when two cars will meet is 9 am + 6.5 h = 3.5 pm.

The time when the second car will overtake the first one is one second more than 3.5 pm

Answer: one second more than 3.5 pm